# LLM Reasoning Engine: Specialized Training for Enhanced Mathematical Reasoning

### v1

@@ Abstract
Large Language Models (LLMs) have shown remarkable performance in various natural language processing tasks but face challenges in mathematical reasoning, where complex problem-solving requires both linguistic understanding and mathematical reasoning skills. Existing approaches to address this challenge often rely on ensemble methods and suffer from the problem of data scarcity in target domains. In this work, we present a novel method to enhance LLMs’ capabilities in mathematical reasoning tasks. Motivated by the need to bridge this gap, our approach incorporates a question paraphrase strategy, which aims at diversifying the linguistic forms of mathematical questions to improve generalization. Additionally, specialized training objectives are employed to guide the model’s learning process, focusing on enhancing its understanding of mathematical concepts and reasoning processes. We conduct experiments on four datasets using different LLMs, and demonstrate the effectiveness of our approach in improving LLMs’ performance on mathematical reasoning tasks. Our findings underscore the significance of our methodology in the advancement of large language models and its potential implications for real-world applications that require mathematical reasoning abilities.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) (Touvron et al., 2023b; Team et al., 2023) have emerged as powerful tools in the field of machine learning, demonstrating remarkable performance in a wide range of downstream benchmarks. Their ability to understand and generate natural language text has revolutionized various applications, from language translation (Costa-jussà et al., 2022; Barrault et al., 2023) to question answering systems (Chowdhery et al., 2023; Chen et al., 2023). Central to their success is their ability to complex reasoning, enabling them to tackle complex problems with impressive accuracy and efficiency.
@@
Some existing methods (Shen et al., 2023; Jiang et al., 2024) attempt to address these challenges by employing ensemble techniques, where multiple LLMs collaboratively solve mathematical problems. However, these methods can still be further improved, particularly in enhancing the performance of individual LLMs and mitigating performance degradation over extended reasoning steps. A major challenge when applying LLMs to complex mathematical problems is error propagation, especially in tasks requiring long reasoning chains. Figure 1 illustrates this issue with an example from the GSM8K dataset (Cobbe et al., 2021) solved by the Llama2-7B model (Touvron et al., 2023b). The problem involves multiple intermediate steps to reach the correct solution. However, the model makes an error in the initial steps, which cascades through subsequent stages of reasoning. This occurs because each reasoning step depends on the accuracy of preceding steps; even a minor initial error can compound, ultimately leading to significant deviations from the correct solution. This phenomenon highlights the limitations of current LLMs in maintaining accuracy in long reasoning sequences. It also underscores the need for improved methodologies to mitigate error propagation and enhance performance in complex mathematical tasks.
@@ 2 Background
The intersection of machine learning and mathematical reasoning has garnered significant attention from researchers seeking to enhance the capabilities of Large Language Models (LLMs) in solving complex mathematical problems. In this section, we review the previous literature exploring various approaches and methodologies to address the challenges posed by mathematical reasoning tasks and highlight the advances made in this domain.
@@ 2.1 Recent advances in LLMs
The evolution of Large Language Models (LLMs) represents a significant milestone in the field of natural language processing (NLP) and machine learning. Over the past decade, advances in neural network architectures, coupled with the availability of large amounts of text data, have driven the development of increasingly complex and capable LLMs. Beginning with seminal works such as Google’s BERT (Bidirectional Encoder Representations from Transformers) (Devlin et al., 2018) and OpenAI’s GPT (Generative Pre-trained Transformer) (Radford et al., 2019), researchers have made substantial progress in enhancing LLMs’ language understanding and generation capabilities (Liu et al., 2019; Raffel et al., 2020). These models leverage transformer architectures and self-attention mechanisms to capture long-range dependencies and contextual information, enabling them to generate coherent and contextually relevant text.
@@ 2.2 LLMs for mathematics
The application of large language models (LLMs) to mathematical reasoning tasks represents a recent and growing area of research within the machine learning community. Although LLMs have demonstrated remarkable proficiency in natural language understanding, their performance in mathematical reasoning tasks has historically lagged behind.
@@
Moreover, novel adaptation techniques, such as question paraphrase methods (Yu et al., 2023) and tailored training objectives (Liu et al., 2023) aimed at enhancing LLMs’ understanding of mathematical reasoning processes, have further contributed to the advancement of LLMs in mathematical applications. These techniques enable LLMs to effectively leverage existing data and develop robust reasoning abilities, thus expanding their utility in mathematical problem solving scenarios. However, these existing methods suffer from the problem of error propagation over long reasoning paths.
@@ 3 Methodology
The motivation for our methodology arises from the observation that while Large Language Models (LLMs) excel in various natural language processing tasks, their performance in mathematical reasoning remains suboptimal. This gap is primarily due to the inherent complexity of mathematical problems, which often requires intricate reasoning and logical deduction (Saxton et al., 2019; Lightman et al., 2023). Furthermore, the limited availability of annotated data in the mathematical domain (Liu et al., 2021; Kumar et al., 2022) presents a significant obstacle to effectively training LLMs for such tasks.
@@ 3.1 Question paraphrase
Question paraphrasing is a crucial technique employed to augment existing data and enhance the model’s ability to generalize across different linguistic forms of mathematical problems. This process involves leveraging the powerful generative capabilities of the GPT-4 model (Brown et al., 2020) to produce diverse paraphrases for each question in the dataset. By generating multiple variations of the same question while preserving its semantic meaning, our aim is to enrich the training data and expose the model to a wider range of linguistic structures and expressions commonly encountered in mathematical problem-solving scenarios.
@@
The GPT-4 model is utilized to generate the paraphrases for each question in the data set. Given a mathematical question, the model generates alternative phrasings that convey the same underlying mathematical concept. This step significantly expands the diversity of the training data by presenting questions in different linguistic forms, including synonyms, rephrasings, and syntactic variations.
Validation via model consistency
@@ 3.3 Training process
The training process of our proposed method consists of several stages designed to enhance the mathematical reasoning capabilities of Large Language Models (LLMs). The process begins with data augmentation through question paraphrasing, where GPT-4 generates various paraphrases for each mathematical question. To ensure data set quality and integrity, we verify the consistency of the paraphrased questions and their corresponding answers, iterating this process to achieve high accuracy.
@@
The final training objective is defined as the weighted sum of $\mathcal{L}_{SFT}$, $\mathcal{L}_{RR}$, and $\mathcal{L}_{MI}$:
@@ 4 Experiments
Mathematical reasoning poses distinct challenges for LLMs, demanding both a deep understanding of natural language and the ability to reason about and manipulate mathematical concepts. Evaluating LLMs’ effectiveness in mathematical reasoning is crucial for advancing state-of-the-art natural language processing and broadening their applicability across diverse problem domains. This study seeks to assess the performance of various Large Language Models (LLMs) on mathematical reasoning tasks through targeted experiments.
@@ 4.1 Datasets
We conducted experiments on four datasets specifically curated to evaluate LLMs’ performance in mathematical reasoning:
@@ 4.2 Base models
We employed four base models for our experiments, each representing a distinct architecture or variant of Large Language Models:
@@
We fine-tuned these base models on the aforementioned datasets using a set of hyperparameters tailored to each model and dataset. The parameters are set as default to the ones used during pre-training. Note that we only experiment with the version of 7B parameters for each model. We adapt LoRA (Hu et al., 2022) to make model fine-tuning more efficient. The rank and alpha are both set as 64. Additionally, we utilized specialized training objectives, such as Rationale Re-ranking (RR) and Mistake Identification (MI), to enhance the models’ understanding of mathematical reasoning. We fine-tune each model with a single Nvidia A100 GPU and the runtime of each experiment is between 2 to 6 hours.
@@ 4.3 Main results
Our experiments yielded notable improvements in the performance of base models across all four datasets. Table 1 presents a summary of the main results. Our proposed special training objectives yield an average performance boost of 4.25% on GSM8K, 2.32% on MATH, 6.21% on GSM_HARD, and 5.15% on SVAMP datasets. Moreover, combining question paraphrase with these objectives further enhances the improvement to 7.32% on GSM8K, 3.63% on MATH, 7.72% on GSM_HARD, and 6.78% on SVAMP. Notably, our methods have a more significant impact on relatively weaker models, with improvements of 9.67% and 9.94% for Llama and Llama2 on GSM8K, compared to 7.14% and 2.52% for Mistral and Mixtral, respectively.
These findings underscore the significance of our approach in enhancing the mathematical reasoning capabilities of Large Language Models. By assessing LLMs’ performance in mathematical reasoning tasks, we contribute to the ongoing efforts to advance the state-of-the-art in natural language processing and pave the way for their application in diverse problem domains requiring mathematical reasoning abilities.
@@ 5 Analysis and discussion
In this section, we delve into a comprehensive analysis and discussion of the experimental results, focusing on the effectiveness of our proposed method in enhancing Large Language Models’ (LLMs) capabilities in mathematical reasoning tasks. We begin by dissecting the model’s performance over varying numbers of reasoning steps, shedding light on the impact of our approach on problem-solving efficiency. Subsequently, we present findings from an ablation study aimed at elucidating the importance of individual components within our proposed methods. Finally, through a series of case studies, we illustrate both the successes and limitations of our approach, providing valuable insights for future research directions.
@@ 5.1 Analysis on reasoning steps
We performed an in-depth analysis of model performance across varying reasoning step counts to evaluate the effectiveness of our proposed method in solving complex mathematical problems. By examining performance at different depths of reasoning, we sought to highlight the impact of our approach on problem-solving efficiency. Figure 3 summarizes the model’s performance on math questions requiring different reasoning steps. Overall, the results indicate that model performance declines as the number of reasoning steps needed to solve the problems increases. Although the improvement is minimal for questions requiring only a few reasoning steps (less than 4 steps), it becomes substantial for questions requiring more extended reasoning chains (4 - 7 steps). Additionally, we note that the models struggle with questions that demand extremely long reasoning paths (more than 8 steps).
Our analysis revealed a clear trend of improvement in model performance as the number of reasoning steps increased. This trend indicates that our proposed method effectively enhances the model’s ability to reason through complex mathematical problems, leading to more accurate solutions. Specifically, we observed substantial performance improvements on problems that involve multiple reasoning steps, which underscores the effectiveness of our approach in addressing complex problem-solving scenarios.
@@ 5.2 Ablation study
To further evaluate the importance of each component in our proposed methods, we performed an ablation study in which we systematically removed individual components and evaluated the model performance. Table 2 presents an ablation study of our proposed method on GSM8K and MATH datasets using Llama2 and Mistral as base models, respectively. Our results demonstrated that all components of our proposed methods play crucial roles in improving model performance in mathematical reasoning tasks.
@@ 5.3 Case study
We present a case study in Figure 4 to illustrate the effectiveness of our proposed method in solving complex mathematical questions. The case study includes a positive example showcasing successful problem-solving outcomes and a negative example highlighting challenges and limitations.
@@
In contrast, the negative example in Figure 4 reveals scenarios where our proposed method encounters limitations or fails to produce satisfactory results. Although the model follows the correct reasoning path, it fails to arrive at the correct solution due to a math calculation error. This example identifies areas for improvement and prompts discussions on potential future research directions, including refining methodologies, leveraging external calculation tools, or exploring math verification approaches to enhance model performance in challenging scenarios.
@@ 6 Conclusion
In this study, we addressed the challenge of enhancing Large Language Models’ (LLMs) capabilities in mathematical reasoning tasks. Our proposed method leverages innovative techniques to improve LLMs’ understanding of mathematical concepts and reasoning processes. Through rigorous experimentation, we demonstrated the effectiveness of our approach in improving LLMs’ performance on various mathematical problems.
Our research contributes to advancing the field of natural language processing by providing a comprehensive approach for enhancing LLMs’ capabilities in mathematical reasoning tasks. By identifying key factors influencing model performance and proposing effective solutions, we bridge the gap between natural language understanding and mathematical reasoning. In conclusion, our study represents a significant step forward in advancing LLMs’ capabilities and expanding their utility in real-world problem-solving scenarios.
@@
While our study demonstrates promising results in enhancing the mathematical reasoning capabilities of LLMs through data enhancement and fine-tuning, several limitations should be acknowledged. First, despite efforts to create a diverse data set through enhancement, inherent biases may persist in the training data. These biases could potentially skew the model’s performance towards certain types of mathematical problems or reasoning patterns. Second, while the fine-tuned model demonstrates improved performance in the test set, its ability to generalize to more complex or novel mathematical problems outside the training distribution remains uncertain. Further investigation is needed to assess the model’s capabilities in tackling advanced mathematical concepts or interdisciplinary problems that significantly differ from the training examples.

### v2

@@ Abstract
Large Language Models (LLMs) have shown remarkable performance in various natural language processing tasks but face challenges in mathematical reasoning, where complex problem-solving requires both linguistic understanding and mathematical reasoning skills. Existing approaches to address this challenge often rely on ensemble methods and suffer from the problem of data scarcity in target domains. In this work, we present a novel method to enhance the capabilities of LLMs in mathematical reasoning tasks. Motivated by the need to bridge this gap, our approach incorporates a question paraphrase strategy, which aims to diversify the linguistic forms of mathematical questions to improve generalization. Additionally, specialized training objectives are employed to guide the model’s learning process, focusing on enhancing its understanding of mathematical concepts and reasoning processes. We conduct experiments on four datasets using different LLMs, and demonstrate the effectiveness of our approach in improving LLMs’ performance on mathematical reasoning tasks. Our findings underscore the significance of our methodology in advancing LLMs and their potential implications for real-world applications that require mathematical reasoning abilities.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) (Touvron et al., 2023b; Team et al., 2023) have emerged as powerful tools in the field of machine learning, demonstrating remarkable performance in a wide range of downstream benchmarks. Their ability to understand and generate natural language text has revolutionized various applications, from language translation (Costa-jussà et al., 2022; Barrault et al., 2023) to question answering systems (Chowdhery et al., 2023; Chen et al., 2023). Central to their success is their ability to perform complex reasoning, enabling them to tackle complex problems with impressive accuracy and efficiency.
@@
Some existing methods (Shen et al., 2023; Jiang et al., 2024) attempt to address these challenges by employing ensemble techniques, where multiple LLMs collaboratively solve mathematical problems. However, these methods can still be further improved, particularly in enhancing the performance of individual LLMs and mitigating performance degradation over extended reasoning steps. A major challenge when applying LLMs to complex mathematical problems is error propagation, especially in tasks requiring long reasoning chains. Figure 1 illustrates this issue with an example from the GSM8K dataset (Cobbe et al., 2021), solved by the Llama2-7B model (Touvron et al., 2023b). The problem involves multiple intermediate steps to reach the correct solution. However, the model makes an error in the initial steps, which cascades through subsequent stages of reasoning. This occurs because each reasoning step depends on the accuracy of preceding steps; even a minor initial error can compound, ultimately leading to significant deviations from the correct solution. This phenomenon highlights the limitations of current LLMs in maintaining accuracy across long reasoning sequences and underscores the need for improved methodologies to mitigate error propagation and enhance performance in mathematical tasks.
@@ 2 Background
The intersection of machine learning and mathematical reasoning has received significant attention from researchers seeking to enhance the capabilities of LLMs in solving complex mathematical problems. In this section, we review the previous literature exploring various approaches and methodologies to address the challenges posed by mathematical reasoning tasks and highlight the advances made in this domain.
@@ 2.1 Recent Advances in LLMs
The evolution of LLMs represents a significant milestone in the field of natural language processing (NLP) and machine learning. Over the past decade, advances in neural network architectures, coupled with the availability of large amounts of text data, have driven the development of increasingly complex and capable LLMs. Beginning with seminal works such as Google’s BERT (Bidirectional Encoder Representations from Transformers) (Devlin et al., 2018) and OpenAI’s GPT (Generative Pre-trained Transformer) (Radford et al., 2019), researchers have made substantial progress in enhancing LLMs’ language understanding and generation capabilities (Liu et al., 2019; Raffel et al., 2020). These models leverage transformer architectures and self-attention mechanisms to capture long-range dependencies and contextual information, enabling them to generate coherent and contextually relevant text.
@@ 2.2 LLMs for mathematics
The application of LLMs to mathematical reasoning tasks represents a recent and growing area of research within the machine learning community. Although LLMs have demonstrated remarkable proficiency in natural language understanding, their performance in mathematical reasoning tasks has historically lagged behind.
@@
Moreover, novel adaptation techniques, such as question paraphrase methods (Yu et al., 2023) and tailored training objectives (Liu et al., 2023) aimed at enhancing LLMs’ understanding of mathematical reasoning processes, have further advanced LLMs in mathematical applications. These techniques enable LLMs to effectively leverage existing data and develop robust reasoning abilities, thus expanding their utility in mathematical problem solving scenarios. However, these existing methods suffer from the problem of error propagation over long reasoning paths.
@@ 3 Methodology
The motivation for our methodology arises from the observation that while LLMs excel in various natural language processing tasks, their performance in mathematical reasoning remains suboptimal. This gap is primarily due to the inherent complexity of mathematical problems, which often requires intricate reasoning and logical deduction (Saxton et al., 2019; Lightman et al., 2023). Furthermore, the limited availability of annotated data in the mathematical domain (Liu et al., 2021; Kumar et al., 2022) presents a significant obstacle to effectively training LLMs for such tasks.
@@ 3.1 Question Paraphrase
Question paraphrasing is a crucial technique employed to augment existing data and enhance the model’s ability to generalize across different linguistic forms of mathematical problems. This process involves leveraging the powerful generative capabilities of the GPT-4 model (Brown et al., 2020) to produce diverse paraphrases for each question in the dataset. By generating multiple variations of the same question while preserving its semantic meaning, we aim to enrich the training data and expose the model to a wider range of linguistic structures and expressions commonly encountered in mathematical problem-solving scenarios.
@@
The GPT-4 model is utilized to generate the paraphrases for each question in the data set. Given a mathematical question, the model generates alternative phrasings that convey the same underlying mathematical concept. This step significantly enhances the diversity of the training data by presenting questions in various linguistic forms, such as synonyms, paraphrases, and syntactic variations.
@@ 3.3 Training Process
The training process of our proposed method consists of several stages designed to enhance the mathematical reasoning capabilities of LLMs. The process begins with data augmentation through question paraphrasing, where GPT-4 generates various paraphrases for each mathematical question. To ensure data set quality and integrity, we verify the consistency of the paraphrased questions and their corresponding answers, iterating this process to achieve high accuracy.
@@
The final training objective is formulated as the weighted sum of $\mathcal{L}_{SFT}$, $\mathcal{L}_{RR}$, and $\mathcal{L}_{MI}$:
@@ 4 Experiments
Mathematical reasoning challenges LLMs, requiring both natural language understanding and mathematical problem-solving skills. Evaluating the effectiveness of LLMs in mathematical reasoning is crucial for advancing state-of-the-art NLP techniques and expanding their applicability across diverse problem domains. This study seeks to assess the performance of various LLMs on mathematical reasoning tasks through targeted experiments.
@@ 4.1 Datasets
We conducted experiments on four datasets specifically designed to assess LLMs’ performance in mathematical reasoning.
@@ 4.2 Base Models
We employed four base models for our experiments, each representing a distinct architecture or variant of LLMs:
@@
We fine-tuned these base models on the aforementioned datasets with hyperparameters tailored to each model and dataset. The parameters are set as default to the ones used during pre-training. Note that we only experiment with the version of 7B parameters for each model. We adapt LoRA (Hu et al., 2022) to make model fine-tuning more efficient. The rank and alpha are both set as 64. Additionally, we utilized specialized training objectives, such as Rationale Re-ranking (RR) and Mistake Identification (MI), to enhance the models’ understanding of mathematical reasoning. We fine-tune each model with a single Nvidia A100 GPU and the runtime of each experiment is between 2 to 6 hours.
@@ 4.3 Main Results
Our experiments resulted in notable performance improvements across all four base models. Table 1 presents a summary of the main results. Our proposed special training objectives yielded an average performance boost of 4.25% on GSM8K, 2.32% on MATH, 6.21% on GSM_HARD, and 5.15% on SVAMP datasets. Moreover, combining question paraphrase with these objectives further enhances the improvement to 7.32% on GSM8K, 3.63% on MATH, 7.72% on GSM_HARD, and 6.78% on SVAMP. Notably, our methods have a more significant impact on relatively weaker models, likely because these models benefit more from structured reasoning guidance. This aligns with findings in previous LLM fine-tuning studies, where weaker models exhibit larger relative improvements when exposed to specialized training objectives.
These findings underscore the empirical effectiveness of our methodology in improving the reasoning efficiency and accuracy of LLMs. By assessing LLMs’ performance in mathematical reasoning tasks, we contribute to the ongoing efforts to advance the state-of-the-art in natural language processing and pave the way for their application in diverse problem domains requiring mathematical reasoning abilities.
@@ 5 Analysis and Discussion
In this section, we delve into a comprehensive analysis and discussion of the experimental results, focusing on the effectiveness of our proposed method in enhancing LLMs’ capabilities in mathematical reasoning tasks. We begin by dissecting the model’s performance over varying numbers of reasoning steps, shedding light on the impact of our approach on problem-solving efficiency. Subsequently, we present findings from an ablation study aimed at elucidating the importance of individual components within our proposed methods. Finally, through a series of case studies, we illustrate both the successes and limitations of our approach, providing valuable insights for future research directions.
@@ 5.1 Analysis on Reasoning Steps
We performed an in-depth analysis of model performance across varying reasoning steps to evaluate the effectiveness of our proposed method in solving complex mathematical problems. By examining performance at different depths of reasoning, we sought to highlight the impact of our approach on problem-solving efficiency. Figure 3 summarizes the model’s performance on math questions requiring different reasoning steps. Overall, the results indicate that model performance declines as the number of reasoning steps needed to solve the problems increases. Although the improvement is minimal for questions requiring only a few reasoning steps (fewer than 4 steps), it becomes substantial for questions requiring longer reasoning chains (4–7 steps). Additionally, we note that the models struggle with questions that demand extremely long reasoning paths (more than 8 steps).
Our analysis revealed a clear trend of improved model performance as the number of reasoning steps increased. This trend indicates that our proposed method effectively enhances the model’s ability to solve complex mathematical problems, leading to more accurate solutions. Specifically, we observed substantial performance improvements on problems that involve multiple reasoning steps, which underscores the effectiveness of our approach in addressing complex problem-solving scenarios.
@@ 5.2 Ablation Study
To further evaluate the importance of each component in our proposed methods, we performed an ablation study in which we systematically removed individual components and evaluated the model performance. Table 2 presents an ablation study of our proposed method on GSM8K and MATH datasets using Llama2 and Mistral as base models, respectively. Our results demonstrated that all components of our proposed methods are integral to improve model performance in mathematical reasoning tasks.
@@ 5.3 Case Study
We present a case study in Figure 4 to demonstrate the effectiveness of our method in solving complex mathematical problems. The case study includes a positive example showcasing successful problem-solving outcomes and a negative example highlighting challenges and limitations.
@@
In contrast, the negative example in Figure 4 reveals scenarios where our proposed method encounters limitations or fails to produce satisfactory results. Although the model follows the correct reasoning path, a common failure mode observed was arithmetic miscalculations despite correct reasoning paths. This suggests that while LLMs grasp mathematical structure, they struggle with precise computation—an issue that could be mitigated by integrating external calculation modules. This example identifies areas for improvement and prompts discussions on potential future research directions, including refining methodologies, leveraging external calculation tools, or exploring math verification approaches to enhance model performance in challenging scenarios.
5.4 Discussion
Data Efficiency and Generalization
Our question paraphrase strategy addresses a fundamental challenge in mathematical reasoning: the scarcity of diverse high-quality training data. By systematically transforming existing questions into linguistically varied forms while preserving their mathematical essence, we achieve improved generalization without requiring additional annotated examples. This approach is particularly valuable for specialized mathematical domains where expert annotation is costly and time-consuming. Our experimental results demonstrate that models trained with paraphrased questions exhibit enhanced robustness to linguistic variations, more closely mirroring the diverse ways in which mathematical problems may be encountered in real-world applications. This linguistic flexibility, combined with our specialized training objectives, enables models to focus on the underlying mathematical structures rather than become overly sensitive to specific phrasings. The data efficiency of our approach makes it particularly applicable to resource-constrained settings and suggests promising directions for self-supervised learning techniques that leverage mathematical invariance across different problem formulations.
Training-focused V.S. Test-time computational methods
In recent years, the Test-Time Computation (TTC) paradigm has gained significant traction as a means to improve the reasoning skills of LLMs, particularly for complex tasks. Although training-focused and TTC methods represent different strategies, they are not mutually exclusive. In fact, their synergistic integration holds significant promise for future advancements in mathematical reasoning for LLMs. The reasoning abilities that TTC techniques leverage are fundamentally learned during the training phase, encompassing both pre-training and fine-tuning. Training on high-quality mathematical data, including examples of step-by-step reasoning, provides the essential knowledge base that enables TTC methods (e.g., CoT (Wei et al., 2022) and Self-Refine (Madaan et al., 2023)) to be effective. Furthermore, training processes can be optimized to produce models that are particularly adept at generating effective reasoning traces, which can then be further explored or verified during test time using TTC techniques. The quality of the underlying model, shaped by training, significantly influences how effectively it can utilize TTC methods.
@@ 6 Conclusion
In this study, we addressed the challenge of enhancing LLMs’ capabilities in mathematical reasoning tasks. Our proposed method leverages innovative techniques to improve LLMs’ understanding of mathematical concepts and reasoning processes. Through rigorous experimentation, we demonstrated the effectiveness of our approach in improving LLMs’ performance on various mathematical problems.
Our research advances the field of natural language processing by providing a comprehensive approach to enhancing LLMs’ capabilities in mathematical reasoning tasks. By identifying key factors influencing model performance and proposing effective solutions, we bridge the gap between natural language understanding and mathematical reasoning. Our study advances LLMs’ reasoning capabilities, paving the way for broader applications in automated theorem proving, mathematical education, and AI-assisted research in scientific domains. Future work could explore integrating symbolic computation with LLMs to further enhance mathematical reliability.
@@
While our study demonstrates promising results in improving the reasoning efficiency and accuracy of LLMs through data enhancement and fine-tuning, several limitations should be acknowledged. First, despite efforts to create a diverse data set through enhancement, inherent biases may persist in the training data. These biases could potentially skew the model’s performance towards certain types of mathematical problems or reasoning patterns. Second, although our fine-tuned model improves test performance, its generalization to out-of-distribution mathematical problems remains uncertain. Future work could explore curriculum learning strategies or hybrid neural-symbolic approaches to mitigate this limitation. Further investigation is needed to assess the model’s capabilities in tackling advanced mathematical concepts or interdisciplinary problems that deviate significantly from the training examples.